OpenStudy (obviouslyhes):

ill medal whoever answers!! Find the total area for the cylinder with the given measurement. r = 2, h = 4 16pi 20pi 24pi

sam (.sam.):

When you open a cylinder, you will get 2 circles and a rectangle. The area of a circle is \(\pi r^2\), we have two circles so \(2 \pi r^2\) and the area of that rectangle is \(2 \pi r h\). Add them up we get \[A=2 \pi r^2 + 2 \pi r h\] Now you have \(r\) and \(h\), can you find the area?

OpenStudy (obviouslyhes):

i got 75.36. is there another step after that to get the final answer?

sam (.sam.):

Nope, it is done.

OpenStudy (obviouslyhes):

but one of the answers has to be 16pi, 20pi, or 24pi.

OpenStudy (mathmale):

I agree with Sam's formula for the total surface area of your cylinder:\[A=2 \pi r^2 + 2 \pi r h\]

OpenStudy (mathmale):

Would you please show your work? It'd then be easier for someone to give you feedback.

TheSmartOne (thesmartone):

Your answers are exact. Plug in r = 2 and h = 4 in to the equation BUT do absolutely nothing to \(\pi\). Because your final answer contains pi, we don't want to get rid of it by doing pi = 3.14 and plugging that in. So plug in r = 2 and h = 4 and simplify what you get without touching pi :)